Timothy has an opportunity to buy a $4,000 par value municipal bond with a coupon rate of

6​% and a maturity of five years. The bond pays interest annually. If Timothy requires a return of 8​%,

what should he pay for the​ bond?

If Timothy requires a return of 8​%,

the amount he should pay for the bond is ??

​(Round to the nearest​ cent.)
Answer #1

Annual coupon=4000*6%=240

Hence price of bond=Annual coupon*Present value of annuity factor(8%,5)+$4000*Present value of discounting factor(8%,5)

=240*3.992710037+$4000*0.680583197

=$3680.58(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=240[1-(1.08)^-5]/0.08

=240*3.992710037

2.Present value of discounting factor=4000/1.08^5

=4000*0.680583197